Session: May 1997 (session 2)
Frequency: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Thu May 29 01:44:19 METDST 1997) Success
No obvious problems

Medicina (T. Venturi at Thu May 29 01:46:58 METDST 1997) Minor failures
The experiment ran smoothly until 21:40 UT. Electrical problems to the engines occurred at that time. The antenna was not able to track the sources until the end of the experiment.

Onsala (A. Polatidis at Thu May 29 10:48:37 METDST 1997) Success
No known problems

Westerbork (ton foley at Thu May 29 11:39:13 METDST 1997) Success
Looks OK

Effelsberg (dgraham at Mon Jun 2 13:47:15 METDST 1997) Success

Jodrell (Alastair Gunn at Thu Jun 5 10:22:17 METDST 1997) Success
No known problems

Cambridge (Alastair Gunn at Thu Jun 5 11:24:17 METDST 1997) Minor failures
Fomatter time out by 10 seconds at start, fixed at 13:08 UT (do not know when this jump occurred). Tape came off at 20:04, restarted FS ready for scan 3 at 20:19 UT (first few minutes missing).

Urumqi (Li Suqin at Fri Jun 6 04:33:43 METDST 1997) Minor failures
It is Cloudy. Clock offset =1.90+- 0.5us. From 14:08:40 to 14:18:20 From15:35:40 to 15:58:40 There are some trouble at antenna. No test were token. The first source was not be tested. So Every souce was tested used the time schedule before it.

Noto (C.Trigilio at Mon Jun 9 15:47:23 METDST 1997) Success
Lost about 8 min from 15:35 UT for cable wrap problem (antenna mouved manually). No other problem.

Simeiz (A.Volvatch/LIG at Wed Jun 18 09:01:06 METDST 1997) Success
No problem reported, tapes shipped on 06.06.97.

Calibration (Fredrik T Rantakyro at Mon Jul 21 11:13:19 METDST 1997) Minor failures
The ANTAB file for the experiment can be found on VLBEER in the MAY97 directory with the name EG17B.ANTAB. The file is ready to be used with the AIPS tasks ANTAB/APCAL. Read the ANTAB file for any specific comments regarding this experiment.